# Aligning Agents like Large Language Models

### v1

@@ Appendix F Reward Model Performances by Jumppad
[…]
Appendix G Alignment of Base Model for Comparison with Fine-Tuned Model
To understand how trajectory diversity affects alignment, we also consider aligning the base agent using the same reward models (trained on the fine-tuned agent). By plotting the successful trajectories that reach the left and right jumppads we see that the base agent (left) has a greater diversity than the fine-tuned agent (right). However, we note that the success rate for reaching the left and right jumppads is lower for the base model (as in Figures 3 and 4), so there are fewer trajectories.
@@ Appendix G Alignment of Base Model for Comparison with Fine-Tuned Model
We now align the base agent to reach the left and right jumppads using the reward models trained on the fine-tuned agent, as shown below. We find that the agent can be aligned more symmetrically with both the left and the right jumppads, although alignment with the left jumppad still has a slightly better performance. However, in comparison to alignment of the fine-tuned agent (Figures 6 and 8), we see that the alignment is much slower, starting at a lower success rate, not increasing as quickly during training, and reaching a lower final success rate. As before, we see the same general trend that higher accuracy reward models (using more preference data) lead to better alignment.
Figure 16: Left jumppad success rate for base agent using fine-tuned agent reward models.
Figure 17: Right jumppad success rate for base agent using fine-tuned agent reward models
Appendix H Final Aligned Agent Jumppad Distributions
@@
Appendix I Limitations and Further Work
@@
This work provides a proof of concept of an approach for training agents to act as desired on a complex 3D environment. For our procedure to be applied in practice by game designers or in other domains such as robotics, the efficiency of this pipeline will be essential.
@@
One limitation of our approach that we discuss in Section 5 is that we utilize synthetic preferences. While this may be a realistic option for many use cases, real human feedback from a game designer will be required in general. As well as potentially introducing noise to the preference labels, this process could be costly in terms of human time, especially for agents on modern console games which must often be run in real-time in order to be rendered. Therefore supervised fine-tuning on relevant behaviours to improve preference labelling efficiency could help to reduce this human time requirement. However, knowledge transfer from training LLMs could again be relevant here. Recent work on providing human feedback for LLMs has used a hybrid form of feedback that combines preference and evaluative feedback, in which responses are grouped into a batch of size $N$ and simultaneously compared on a preference scale of size $P$, which enables larger numbers of comparisons to be extracted from a given number of responses, improving the time efficiency of providing feedback. For example, InstructGPT (Ouyang et al., 2022) uses $4\leq N\leq 9$, $P=7$ while Llama2 (Touvron et al., 2023) uses $N=2$, $P=4$. As $N$ and $P$ increase, more information can be extracted from the provided human feedback. For example, with $N=5$, $P=5$ and assuming no category duplication (no trajectories are considered equal), (${5\choose 2}=10$ comparisons can be extracted, which is equivalent to 2 bits of information per trajectory watched, compared to just 0.5 bits per trajectory for default pairwise comparisons with $N=2$, $P=2$. This results in a 4×4$\times$4 × improvement in feedback efficiency for human labellers, at the cost of potential label noise due to the additional mental overhead required. Similar strategies could be applied to providing feedback to agents to make the process more time efficient.
@@
Another limitation we discuss is that running agents online in modern console games can be challenging. While our work demonstrated surprising efficiency for fine-tuning pre-trained agents with RL, for more complex behaviours and tasks this may become prohibitively expensive. Fortunately, developments in RLHF for language models can also be utilized here. We demonstrated in Section 3 that an intial step of preference fine-tuning could improve the efficiency of alignment, and explained that this corresponds to Reinforced Self-Training (ReST) (Gulcehre et al., 2023) with a single iteration. However, the full procedure using multiple iterations could be used to further improve the sample efficiency of aligning with preferences while maintaining the benefits of online exploration. Other recent work has investigated aligning models completely offline. Direct Preference Optimisation (Rafailov et al., 2023) provides an approach for optimising a model to align with preferences without the need for a reward model or online training, while (Hu et al., 2023) uses offline reinforcement learning with pre-generated samples and rewards. Additionally, efficient fine-tuning strategies such as LoRA (Hu et al., 2021) could be used to reduce the hardware requirements for the fine-tuning of a large base model for the individual applications of game designers. These new approaches provide promising directions for further work in the context of agents to improve the efficiency or potentially even remove the need for online training completely. However, the extent to which active online learning is required for true generality is still an open question.

### v2

@@ Appendix F Reward Model Performances by Jumppad
Both jumppads show the same trend for the agent encoder, reaching $\sim 90\%$ performance for $100$ comparisons and $100\%$ performance by $500\text{k}$ comparisons. The random encoder is high variance, but generally increases in performance with training comparisons for the left jumppad, while decreasing in performance for the right jumppad. This surprising trend for the right jumppad could be because the reward model begins to overfit to the imbalanced trajectories discussed in Section 3.6.2 and shown in Figure 15. In other words, the reward model misinterprets the preferences and instead learns to prefer shorter trajectories that begin at the bottom left spawn point (from which 84% of the successful right-going trajectories arise) leading to the $\sim 20\%$ accuracy that we see for higher numbers of comparisons. The agent encoder based reward model is better able to distinguish these trajectories and infer the true underlying preference that better fits the data.
Appendix G Motivation for the Use of REINFORCE for Agent Alignment
Our proof-of-concept uses REINFORCE (williams_simple_1992) for online alignment of our agent, as discussed in Section 3.6. While much of the RLHF literature uses PPO (schulman_proximal_2017), we note that recent work on LLMs has found PPO to be unnecessarily complex for RLHF (ahmadian_back_2024). This is due to the fact that in the standard RLHF framework, the reward is only provided at the end of the trajectory, meaning that we generally just want to reinforce entire trajectories rather than attempt credit assignment at each timestep using an advantage function, as is typically done with PPO. Additionally, ahmadian_back_2024 found that importance clipping rarely occurs in practice in RLHF, since the final aligned policy is kept close to the initial fine-tuned policy. This is reflected by the use of REINFORCE in recent state-of-the-art LLMs (google_deepmind_gemma_2024). As a result, we use REINFORCE for simplicity in our proof-of-concept, but note that more complex tasks or feedback mechanisms may warrant the additional complexity of more modern reinforcement algorithms such as PPO.
Appendix H Investigation into Discrepancy in Aligning Agent to Reach Right Jumppad
H.1 Understanding Discrepancy between Left and Right Jumppad Trajectories
Since the task of aligning our agent left or right is seemingly equivalent, we would expect to be able to align the agent similarly. Surprisingly however, we found that our fine-tuned agent did not align as quickly towards the right jumppad as the left, only reaching around a $\sim 40\%$ success rate (in the absence of preference fine-tuning) after one day of training (as shown in Figure 9) compared to around $\sim 90\%$ for the left jumppad (as shown in Figure 8). In order to investigate this discrepancy, we analyzed the rollouts of the fine-tuned agent, shown in Figure 15.
We see that the trajectories that reach the left jumppad start relatively evenly from the 4 spawn locations. However, of the trajectories that successfully reach the right jumppad, only 1% originated from the spawns on the right hand side. Therefore it is unsurprising that it is more difficult for the agent to learn to go right, given that in half the episodes (from the right spawn locations) the agent rarely navigates to the right jumppad to receive reward to reinforce its behavior. To confirm this hypothesis, we instead attempt to align the base agent (before fine-tuning).
H.2 Alignment of Base Model for Comparison with Fine-Tuned Model
To understand how trajectory diversity affects alignment, we now consider aligning the base agent using the same reward models (trained on the fine-tuned agent). By plotting the successful trajectories that reach the left and right jumppads we see that the base agent (left) has a greater diversity than the fine-tuned agent (right). However, we note that the success rate for reaching the left and right jumppads is lower for the base model (as in Figures 3 and 4), so there are fewer trajectories.
@@ H.2 Alignment of Base Model for Comparison with Fine-Tuned Model
We now align the base agent to reach the left and right jumppads using the reward models trained on the fine-tuned agent, as shown below. We find that the agent can be aligned more symmetrically with both the left and the right jumppads, although alignment with the left jumppad still has a slightly better performance. However, in comparison to alignment of the fine-tuned agent (Figures 8 and 9), we see that the alignment is much slower, starting at a lower success rate, not increasing as quickly during training, and reaching a lower final success rate. As before, we see the same general trend that higher accuracy reward models (using more preference data) lead to better alignment.
Figure 17: Left jumppad success rate for base agent using
fine-tuned agent reward models.
Figure 18: Right jumppad success rate for base agent using
fine-tuned agent reward models
Appendix I Final Aligned Agent Jumppad Distributions
@@
Appendix J Limitations and Further Work
@@
Our work provides a proof-of-concept of an approach for training embodied agents to act as desired in complex 3D environments. For this procedure to be applied in practice in video games and other domains such as robotics, the efficiency of this pipeline will be essential.
One limitation of our approach is that we assume access to a large amount of human data on the environment of interest to pre-train a base model. In practice however, when designing a game, large amounts of human data may not yet be available, and even for existing games it may not be possible to access. A potential solution may be to use a ‘gaming foundation model’ trained on similar games (with pixel input and action controller output). Equivalently in robotics, the collection of general cross-objective and embodiment datasets is becoming more important for training general robotic agents (embodimentcollaboration2024openxembodimentroboticlearning). We believe the potential transfer of models pre-trained across environments will be an exciting direction for future work.
Investigating developments in LLM architectures and pre-training could also be of interest for further work. For example, the context window used by the transformer $H$ corresponds to the memory of the agent, so developments in increasing or augmenting the context window in LLMs such as Longformer attention (beltagy_longformer_2020) or Retrieval-Augmented Generation (RAG) (lewis_retrieval-augmented_2021), could similarly be incorporated to increase the memory of agents trained with this approach, helping to address the issues with partial observability in complex 3D environments.
@@
Another limitation of our proof-of-concept that we discuss in Section 3.4 is that we utilize synthetic preferences. While this may be a realistic option for many use cases, potentially using VLMs-as-a-judge (zheng_judging_2023; li_llms-as-judges_2024), real human feedback may be required in general. As well as potentially introducing noise to the preference labels, this process can be costly in human time, especially for agents on modern console games which must often be run in real-time in order to be rendered. Therefore supervised fine-tuning on relevant behaviors to improve preference labeling efficiency could help to reduce this human time requirement. However, knowledge transfer from training LLMs could again be relevant here. Recent work on providing human feedback for LLMs has used a hybrid form of feedback that combines preference and evaluative feedback, in which responses are grouped into a batch of size $N$ and simultaneously compared on a preference scale of size $P$, which enables larger numbers of comparisons to be extracted from a given number of responses, improving the time efficiency of providing feedback. For example, InstructGPT (ouyang_training_2022) uses $4\leq N\leq 9$, $P=7$ while Llama2 (touvron_llama_2023) uses $N=2$, $P=4$. As $N$ and $P$ increase, more information can be extracted from the provided human feedback. For example, with $N=5$, $P=5$ and assuming no category duplication (no trajectories are considered equal), ${5\choose 2}=10$ comparisons can be extracted, which is equivalent to 2 bits of information per trajectory watched, compared to just 0.5 bits per trajectory for default pairwise comparisons with $N=2$, $P=2$. This results in a $4\times$ improvement in feedback efficiency for human labellers, at the cost of potential label noise due to the additional mental overhead required. Similar strategies could be applied to providing feedback to agents to make the process more time efficient.
@@
A final limitation we consider in our proof-of-concept is the ability to run agents online, since embodied agents are much more challenging to run online than LLMs, even in the context of video games. For more complex behaviors and tasks this may become prohibitively expensive. Fortunately, developments in RLHF for language models can also be utilized here. We demonstrated in Section 3.6 that an initial step of preference fine-tuning, corresponding to Reinforced Self-Training (ReST) (gulcehre_reinforced_2023) with a single iteration, can improve alignment efficiency. However, the full procedure using multiple iterations could be used to further improve the sample efficiency of aligning with preferences while maintaining the benefits of online exploration. Other recent work has investigated aligning models completely offline. Direct Preference Optimization (rafailov_direct_2023) provides an approach for optimizing a model to align with preferences without the need for a reward model or online training, while (hu_aligning_2023) uses offline reinforcement learning with pre-generated samples and rewards. Additionally, efficient fine-tuning strategies such as LoRA (hu_lora_2021) could be used to reduce the hardware requirements for the fine-tuning of a large base model for individual use-cases. These new approaches provide promising directions for further work in the context of agents to improve the efficiency or potentially even remove the need for online training completely. However, the extent to which active online learning is required for true generality is still an open question (ostrovski_difficulty_2021) as we discuss in Appendix D.
In summary, we believe there is significant potential for transfer from recent developments in large language models to agents, that could help to address many long-standing limitations and open-problems for decision-making agents. We believe these will open up exciting new research directions in the coming years.